OpenStudy (anonymous):

Please Help Ill give medals & fans . Describe a real life situation that can be represented as a system of inequalities.

OpenStudy (anonymous):

Erm if one store sells something cheaper (X) than the other(y). x<y

OpenStudy (anonymous):

Here i'll make one for you. x=The cost of a Burger y=The cost of a Drink Miko goes to a fast food returant and buys 6 burgers and two drinks for a total of $ 14 before tax. Based on Miko's purchase we know 6x+2y=14 To solve I can find the value of y in terms of x in Miko's equation. 6x+2y=14 subtract 6x on both sides 2y=14-6x divide by 2 on both sides y=7-3x Since it is a System there are 2 scenarios that go with each other. Sarah goes to a fast food resturant and buys 4 burgers and 4 drinks for a total of $ 16 before tax. Based on Sarah's purchase we know 4x+4y=16 Now plug the value from jacks equation in place of y in Sarah's equation 4x+4y=16 so 4x+4(7-3x)=16 4x+28-12x=16 combine like terms 28-8x=16 subtract 28 on both sides -8x=-12 Divide by -8 on both sides x=1.5 So the cost of a burger is $1.50 We also showed that y=7-3x Since x=1.5 we can substitute $1.50 for x y=7-3(1.5) y=7-4.5 y=2.5 So the cost of a Drink is $2.50 Burgers: %1.50 Drinks: $2.50 Since a Drink and a Burger cost $1.50 plus $2.50 or 4 dollars John would not have enough money in question B.

OpenStudy (anonymous):

Oh boy that was a lot -_-

OpenStudy (anonymous):

Any Questions?

OpenStudy (anonymous):

Lol that is A lot thank you so much! @Alexandra675

OpenStudy (anonymous):

any questions though?

OpenStudy (anonymous):

no @Alexandra675

OpenStudy (anonymous):

ok and You're Welcome :)